# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 7.2 Tree-like single-coalition-first neighborhood models
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
@@
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ from $r$ to $s$.
@@
    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $a\in\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$;
@@
    there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
@@
Assume (1). Let $r$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $a$-history from $r$ to $s$.
Let $s$ be a state of $\mathtt{S}\textbf{-}\mathtt{NM}$ and $\mathrm{C}\subseteq\mathtt{AG}$.
  • Assume $s=r$. Then, $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{m},s_{m})$ where $s_{0}=s_{m}=r$ is a $\mathrm{C}$-history from $r$ to $r$ which is different from $r$. For every $1\leq i\leq m$, let $Z_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $Z_{i}\subseteq\mathrm{Y}_{i}$ and $s_{1}\in Z_{i}$. Then, $(s_{0},Z_{1},s_{1},\dots,Z_{m},s_{m})$ is an $a$-history from $r$ to $r$ which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
@@
    Assume $\mathrm{C}=\emptyset$. Let $a\in\mathtt{AG}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$. It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$ is a unique $a$-history from $r$ to $s$.
    Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ be the $a$-history from $r$ to $s$.
    For every $1\leq i\leq n$, let $Z_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq Z_{i}$. Then, $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
    Let $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ be a $\mathrm{C}$-history from $r$ to $s$. We claim $(s_{0},Z_{1},s_{1},\dots,Z_{n},s_{n})$ and $(t_{0},D_{1},t_{1},\dots,D_{m},t_{m})$ are identical.
    For every $1\leq i\leq m$, let $E_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $D_{i}\subseteq E_{i}$. Then, $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$ is a $a$-history from $r$ to $s$.
    Assume $n\neq m$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
    Assume $s_{i}\neq t_{i}$ for some $1\leq i\leq n$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
    Assume $Z_{i}\neq D_{i}$ for some $1\leq i\leq n$. Then, $Z_{i}\cap D_{i}=\emptyset$. Then, $s_{i}\neq t_{i}$. Then, $(s_{0},\mathrm{Y}_{1},s_{1},\dots,$ $\mathrm{Y}_{n},s_{n})$ is different from $(t_{0},E_{1},t_{1},\dots,E_{m},t_{m})$. We have a contradiction.
@@
It easy to see that (1) is a special case of (2).
By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which are respectively concerned with all agents, all coalitions, and the grand coalition. It would be expected that the following condition is equivalent to the two conditions in Theorem 17: there is a state $r$ of $\mathtt{S}\textbf{-}\mathtt{NM}$ such that for every $\mathtt{AG}\subseteq\mathtt{AG}$, for every state $s$ of $\mathtt{S}\textbf{-}\mathtt{NM}$, there is an unique $\mathtt{AG}$-history from $r$ to $s$. In fact, this is not the case. The counterexample given after Theorem 11 also works here.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model. By Theorem 4, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model. Then, it is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model. ∎
@@
Every tree-like grand-coalition-first neighborhood model $z$-represents a tree-like single-coalition-first action model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model.
Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 8. As shown in that proof, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$.
It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.
From Theorem 18, Theorem 19, and Theorem 7, the following result follows:
Theorem 20 (Representation of tree-like grand-coalition-first action models by tree-like single-coalition-first neighborhood models).
@@ 8 Each of those eight coalition logics is determined by these six kinds of models, too
For every finite nonempty sequence $\chi$, we use $\mathsf{last}(\chi)$ to denote the last element of $\chi$.
@@
Let $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model.
There is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ such that for all $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
@@
    • $\mathtt{ST}_{0}^{\prime}=\{w\}$;
    • $\mathtt{ST}_{k+1}^{\prime}=\{\chi-\sigma_{\mathtt{AG}}-u\mid\chi\in\mathtt{ST}_{k}^{\prime},\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}},u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
  • for every $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$:
    $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})=\{y\in\mathtt{ST}^{\prime}\mid y=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\}$.
  • for every $\chi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\chi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\mathsf{last}(\chi))\}$.
Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is what we want.

### v2

@@ 6.2 Tree-like single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
For every $\mathrm{C}\subseteq\mathtt{AG}$, a finite sequence $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$, where every $s_{i}$ is a state and every $\mathrm{Y}_{i}$ is a set of states, is called a $\mathrm{C}$-history from $s_{0}$ to $s_{n}$ if for every $i$ such that $0\leq i<n$, $\mathrm{Y}_{i+1}\in\mathtt{nei}_{\mathrm{C}}(s_{i})$ and $s_{i+1}\in\mathrm{Y}_{i+1}$.
@@
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
The following result is easy to verify and we skip its proof.
Fact 6.
Every tree-like single-coalition-first neighborhood model is a clear single-coalition-first neighborhood model, but not vice versa.
The following result offers an alternative definition of tree-like single-coalition-first neighborhood models.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
@@
Assume (1). Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is a unique $\mathrm{C}$-history from $r$ to $s$.
Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume there is a different $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$. Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
Assume $\mathrm{C}=\emptyset$. Then for every $i$ such that $1\leq i\leq n$, $\mathrm{Y}_{i}=\mathtt{suc}(s_{i-1})$. Let $a\in\mathtt{AG}$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$ for every $s\in\mathtt{ST}$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{m},s_{m})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
@@
Assume $\mathrm{C}=\emptyset$. Let $a\in\mathtt{AG}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. It is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),$ $s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$. Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\mathtt{suc}(t_{0}),t_{1},\dots,\mathtt{suc}(t_{m-1}),t_{m})$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathtt{suc}(t_{i-1})$ and $t_{i}\in\mathrm{Z}_{i}$. Then, $(t_{0},\mathrm{Z}_{1},t_{1},\dots,\mathrm{Z}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for all $0\leq i\leq n$. Then, $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$.
It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\mathrm{U}_{1},t_{1},\dots,\mathrm{U}_{m},t_{m})$. For every $1\leq i\leq m$, let $\mathrm{V}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{U}_{i}\subseteq\mathrm{V}_{i}$. Then, $(t_{0},\mathrm{V}_{1},t_{1},\dots,\mathrm{V}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for every $0\leq i\leq n$. Assume $\mathrm{Z}_{i}\neq\mathrm{U}_{i}$ for some $1\leq i\leq n$. Note $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood model. Then, $\mathrm{Z}_{i}\cap\mathrm{U}_{i}=\emptyset$. Then, $s_{i}\neq t_{i}$, which is impossible. Then, $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$.
@@
(1) is a special case of (2).
By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition is equivalent to the two conditions in Theorem 17: there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$. In fact, this is not the case. The example given after Theorem 11 also works here.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model. By Theorems 16 and 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model. ∎
@@
Every tree-like single-coalition-first neighborhood model $z$-represents a tree-like grand-coalition-first action model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown there, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.
The following result follows from Theorems 18, 19, and 7.
Theorem 20 (Representation of the class of tree-like grand-coalition-first action $\mathtt{X}$-models by the class of tree-like single-coalition-first neighborhood $\mathtt{X}$-models).
@@ 7 Each of those eight coalition logics is determined by these six kinds of models, too
For every $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, there is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime})$ such that for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime}\Vdash\phi$.
@@
Lert $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
In the following, for every finite nonempty sequence $\pi$, we use $\pi^{l}$ to denote the last element of $\pi$.
@@
    • $\mathtt{ST}_{0}^{\prime}=\{s\}$;
    • $\mathtt{ST}_{k+1}^{\prime}=\{\pi-\sigma_{\mathtt{AG}}-u\mid\pi\in\mathtt{ST}_{k}^{\prime},\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}},\text{and }u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\}$.
  • for every $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})=\{\pi-\sigma_{\mathtt{AG}}-u\mid u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\}$.
  • for every $\pi\in\mathtt{ST}^{\prime}$, $\mathsf{label}^{\prime}(\pi)=\{p\in\mathtt{AP}|$ $p\in\mathsf{label}(\pi^{l})\}$.
Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$ and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is what we want.
We first make two claims:
  1. (1)
    for every $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$, if $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, then $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$;
  2. (2)
    for every $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, then there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
First, we show the claim (1). Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$. Assume $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. Note $\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\subseteq\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $u\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, that is, $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
Second, we show the claim (2). Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $t\in\mathtt{ST}$. Assume $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $t\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$. Note $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})\subseteq\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.